# Asymptotic Analysis of High Order IMEX-RK Methods for ES-BGK Model at Navier-Stokes level

### v2

@@ 3 Asymptotic Properties of the IMEX RK Schemes
$U^{n+1}=U^{n}-\Delta t\tilde{b}^{\top}\nabla_{x}F({\bf U})-\varepsilon\Delta t\tilde{b}^{\top}\nabla_{x}H({\bf U}),$ (3.52)
with
$H({\bf U})=\left(\begin{array}[]{c}0\\ \rho{\bf\Theta_{1}}\\ {\bf\mathbb{Q}_{1}}+\rho{\bf\Theta_{1}}{\bf u}\end{array}\right).$
Now we evaluate the quantities $\rho{\bf\Theta_{1}}$ and $\bf\mathbb{Q}_{1}$}. To achieve this, we expand the anisotropic Gaussian $\mathcal{G}$ with respect to $\varepsilon$, i.e.,
@@
$\frac{\mathcal{M}[{{\bf G}}]-\mathcal{\mathcal{M}}[f^{n}]{\bf e}}{\Delta t}+\tilde{A}\,\left(L(\mathcal{M}[{\bf G}])\right)=\tilde{A}\mathcal{M}[{\bf G}]\left(A(V):\frac{\sigma({\bf u})}{2}+2B(V)\cdot\nabla_{x}\sqrt{{\bf T}}\right)+\mathcal{O}(\varepsilon),$ (3.56)
with ${\bf u}=(u^{(1)},...,u^{(s)})^{\top}$, ${\bf T}=(T^{(1)},...,T^{(s)})^{\top}$ and
$V=\frac{v-{\bf u}}{\sqrt{T}},\quad A(V)=V\otimes V-\frac{1}{d_{v}}|V|^{2}Id,\quad B(V)=\frac{1}{2}V(|V|^{2}-(d+2)),$ (3.57)
and $\bf G$} denotes the vector of internal stage for the equation (A.115). Integrating both sides of (3.56) macroscopic variables are associated to $\bf G$} are obtained as follows:
$\begin{split}{\bf U}_{\bf G}&=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U}_{\bf G})+\mathcal{O}(\varepsilon\Delta t)\cr&=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U})+\mathcal{O}(\varepsilon\Delta t),\end{split}$ (3.58)
where ${\bf U}_{\bf G}$ is the macroscopic quantities associated to the stage values $\bf G$} and $F({\bf U})=F({\bf U}_{\bf G})$ follows from the fact that first stage values of $\bf F$} and $\bf G$} are the same $f^{n}$. This together with (3.51) gives $\|{\bf U}-{\bf U}_{\bf G}\|_{\infty}=\mathcal{O}(\varepsilon\Delta t)$, and Lemma A.1 further implies that $\mathcal{M}[{\bf G}]=\mathcal{M}[{\bf F}]+\mathcal{O}(\varepsilon\Delta t)$. Inserting this into (3.55), we get
$\begin{split}\bar{\tau}({\bf f}_{1}-{\bf g})&=-A^{-1}\left(\displaystyle\frac{\mathcal{M}[{\bf G}]-\mathcal{M}[f^{n}]{\bf e}}{\Delta t}+\tilde{A}\,L(\mathcal{M}[{\bf G}])+\mathcal{O}(\varepsilon)+\mathcal{O}(\varepsilon\Delta t)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)\\ &=-A^{-1}\left(\tilde{A}\mathcal{M}[{\bf G}]\left(A(V):\frac{\sigma({\bf u})}{2}+2B(V)\cdot\nabla_{x}\sqrt{{\bf T}}\right)+\mathcal{O}(\varepsilon)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)+\mathcal{O}(\varepsilon\Delta t)\end{split}$ (3.59)
@@
$\bar{\tau}({\bf f}_{1}-{\bf g})$ $=-A^{-1}\left(\tilde{A}\mathcal{M}[{\bf F}]\left(A(V):\frac{\sigma({\bf u})}{2}+2B(V)\cdot\nabla_{x}\sqrt{{\bf T}}\right)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)+\mathcal{O}(\varepsilon)+\mathcal{O}(\varepsilon\Delta t)$
@@
${\bf f}_{1}={\bf g}-{\bar{\tau}}^{-1}A^{-1}\left(\tilde{A}\mathcal{M}[{\bf F}]\left(A(V):\frac{\sigma({\bf u})}{2}+2B(V)\cdot\nabla_{x}\sqrt{{\bf T}}\right)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)+\mathcal{O}(\varepsilon),$ (3.60)
Multiplying by $v\phi$ both side in (3.60) and taking the integration over $v$, we get
$\left\langle v\phi{\bf f}_{1}\right\rangle=\left\langle v\phi{\bf g}\right\rangle-{\bar{\tau}}^{-1}A^{-1}\tilde{A}\left\langle v\phi\left(\mathcal{M}[{\bf F}]\left(A(V):\frac{\sigma({\bf u})}{2}+2B(V)\cdot\nabla_{x}\sqrt{{\bf T}}\right)\right)\right\rangle+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right),$ (3.61)
that is, (Appendix A, see (A.120)),
$H({\bf U})=\mathcal{G}({\bf U})-{\bar{\tau}}^{-1}A^{-1}\tilde{A}\mathcal{F}({\bf U})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right),$ (3.62)
with $H({\bf U})=(H(U^{(1)}),...,H(U^{(s)}))^{\top}$, $\mathcal{G}({\bf U})=(\mathcal{G}(U^{(1)}),...,\mathcal{G}(U^{(s)}))^{\top}$ and $\mathcal{F}({\bf U})=(\mathcal{F}(U^{(1)}),...,\mathcal{F}(U^{(s)}))^{\top}$ obtained by (A.116), (A.117) and (A.118). Thus, from (3.62), the approximations of stress tensor and heat flux are given by
$\rho{\bf\Theta_{1}}=-\bar{\mu}\sigma({\bf u})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right),$
${\bf q}=\mathbb{Q}_{1}=-\bar{\kappa}\nabla_{x}{\bf T}+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right),$
where $\sigma({\bf u})=\nabla_{x}{\bf u}+(\nabla_{x}{\bf u})^{\top}-\frac{2}{d_{v}}\nabla_{x}\cdot{\bf u}Id.$ Here we defined the viscosity and thermal conductivity $s\times s$ matrices ${\bar{\mu}}=(\mu_{ij})$ and ${\bar{\kappa}}=(\kappa_{ij})$ defined by
${\bar{\mu}}=\frac{1}{(1-\nu)}{\bar{\tau}}^{-1}(A^{-1}\tilde{A})\,\bar{p}e^{\top},\quad{\bar{\kappa}}=\frac{d_{v}+2}{2}\,{\bar{\tau}}^{-1}(A^{-1}\tilde{A})\bar{p}e^{\top},$ (3.63)
with $\bar{p}=(p^{(1)},p^{(2)},...,p^{(s)})^{\top}$. This implies that
$\bar{p}e^{\top}=(A^{-1}\tilde{A})^{-1}{\bar{\tau}}(1-\nu){\bar{\mu}}=(A^{-1}\tilde{A})^{-1}{\bar{\tau}}\frac{2}{d_{v}+2}{\bar{\kappa}}$
@@
This form is consistent to the Prandtl number of the ES-BGK model.
@@
${\bf U}=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U})+\varepsilon\Delta t\tilde{A}\nabla_{x}S({\bf U})+\mathcal{O}(\varepsilon^{2}),$ (3.64)
and
$U^{n+1}=U^{n}-\Delta t\tilde{b}^{\top}\nabla_{x}F({\bf U})+\varepsilon\Delta t\tilde{b}^{\top}\nabla_{x}S({\bf U})+\mathcal{O}(\varepsilon^{2}),$ (3.65)
for the numerical solution, where
$S({\bf U})=\left(\begin{array}[]{c}0,\\ \bar{\mu}\sigma({\bf u}),\\ \bar{\mu}\sigma({\bf u}){\bf u}+{\bf q}\end{array}\right).$ (3.66)
We observe that a classical explicit RK method of order $p$ applied to the CNS equations (1.21):
$\partial_{t}U+\nabla_{x}\cdot F(U)=\varepsilon\nabla_{x}\cdot S(U)$
is given by
${\bf U}=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U})+\varepsilon\Delta t\tilde{A}\nabla_{x}S({\bf U}),$ (3.67a)
$U^{n+1}=U^{n}-\Delta t\tilde{b}^{\top}\nabla_{x}F({\bf U})+\varepsilon\Delta t\tilde{b}^{\top}\nabla_{x}S({\bf U}).$ (3.67b)
Now, assuming $u(t)$ the true solution of CNS, the local truncation error is $\mathcal{O}(\Delta t^{p})$, i.e.,
$\begin{split}\mathcal{O}(\Delta t^{p})&=\frac{u(t_{n+1})-u(t_{n})}{\Delta t}+\sum_{i=1}^{s}\tilde{b}_{i}\left(\nabla_{x}F(u^{(i)})-\varepsilon\nabla_{x}S(u^{(i)})\right)\end{split}$ (3.68)
@@
${u}^{(i)}=u(t_{n})-\Delta t\sum_{j=1}^{k}\tilde{a}_{jk}\left(F(u^{(j)})+\varepsilon\nabla_{x}\cdot S(u^{(j)})\right).$
Now, we consider the local trunction of our method (3.65):
$\begin{split}L.T.E.&=\frac{u(t_{n+1})-u(t_{n})}{\Delta t}+\sum_{i=1}^{s}\tilde{b}_{i}\left(\nabla_{x}F(u^{(i)})-\varepsilon\nabla_{x}S(u^{(i)})\right)+\mathcal{O}\left(\frac{\varepsilon^{2}}{\Delta t}\right).\end{split}$ (3.69)
This combined with (3.68), gives
$L.T.E.=\mathcal{O}(\Delta t^{p})+\mathcal{O}\left(\frac{\varepsilon^{2}}{\Delta t}\right)$ (3.70)
Theorem 3.6 implies that, to accurately capture the Navier-Stokes equation, it is necessary that the local truncation error satisfies $L.T.E=o(\varepsilon)$, which holds true if $\Delta t^{p}+\frac{\varepsilon^{2}}{\Delta t}={o}(\varepsilon)$.
In the stiff case, i.e., $\Delta t\gg\varepsilon$, the phenomenon of order reduction may occur, particularly in the worst case when $\Delta t\approx\varepsilon$ (intermediate regime of $\varepsilon$). In this scenario, the classical order of the method can drop off, leading to a loss of accuracy in IMEX-RK schemes within this regime.
In (3.70), in the worse case scenario $\Delta t\approx\varepsilon$ the maximum error is $\mathcal{O}(\Delta t)$ for $p>1$. For instance, this result states that the a third-order schemes $p=3$ may exhibit only first-order accuracy in the worst case. However, this is a lower bound, and in practice, better behavior is often observed (see numerical results).
This phenomenon was investigated in detail in [7, 9], where the authors considered a prototype hyperbolic relaxation system and performed an asymptotic expansion up to $\mathcal{O}(\varepsilon)$ for IMEX-RK methods of type I. They showed that, under additional order conditions, these schemes effectively reduce to explicit Runge-Kutta methods at the $\mathcal{O}(\varepsilon)$ level. In particular, these methods were carefully designed to accurately match the $\mathcal{O}(\varepsilon)$ terms up to a desired order $p$, ensuring that effectively behave as explicit Runge-Kutta schemes of order $p$ for the convection-diffusion equation when $\varepsilon$ is small but not negligible. However, constructing high-order IMEX-RK schemes of type I requires additional stages to satisfy extra-order conditions. In this work, we do not consider high-order IMEX-RK schemes of type I that satisfy these additional conditions introduced in [7], as their construction is particularly challenging. Instead, we focus on type II schemes, which are easier to construct. In Section 4, we compare these methods with other IMEX-RK schemes of type I or II existing in literature, and we will show that the later schemes that do not satisfy the extra order conditions may suffer from order reduction in the intermediate regime of $\varepsilon$.
Finally, considering the results in [7, 9], we state the following.
@@
Setting $\tau=1$, in (1.7), the numerical scheme (3.64)-(3.65) becomes
$\begin{split}{\bf U}&=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U})+\varepsilon\Delta tB\nabla_{x}S({\bf U}),\\[5.69054pt] U^{n+1}&=U^{n}-\Delta t\tilde{b}^{\top}\nabla_{x}F({\bf U})+\varepsilon\Delta t\omega^{\top}\nabla_{x}S({\bf U}),\end{split}$ (3.71)
which is an explicit-type RK method for CNS equation (1.21) based on the coefficients matrices $\tilde{A},B$ and the weights $\tilde{b},w$, with
$B=\tilde{A}A^{-1}\tilde{A},\quad\omega^{\top}=\tilde{b}^{\top}A^{-1}\tilde{A}$,
[…]
In this case, viscosity and thermal conductivity $\bar{\mu}$} and $\bar{\kappa}$} are defined by
${\bar{\mu}}=\frac{1}{(1-\nu)}\bar{p},\quad{\bar{\kappa}}=\frac{d_{v}+2}{2}\bar{p},$ (3.72)
with $\bar{p}=(p^{(1)},p^{(2)},...,p^{(s)})^{\top}$.
@@
IMEX-RK scheme of type II. In this section we analyze the asymptotic behavior of type II IMEX-RK schemes. As type I scheme, at this stage, we do not need to assume that the scheme satisfies the GSA condition.
@@
For small values of $\varepsilon$ with $\Delta t^{p}+\frac{\varepsilon^{2}}{\Delta t}={o}(\varepsilon)$, and $p$, the order of the explicit RK scheme, the IMEX-RK scheme of type II satisfying
$\hat{\tilde{A}}\hat{A}^{-1}({a}-\tilde{a})=0,\quad\hat{\tilde{b}}^{\top}\hat{A}^{-1}({a}-\tilde{a})=0,$ (3.73)

### v3

@@ 3 Asymptotic Properties of the IMEX RK Schemes
$U^{n+1}=U^{n}-\Delta t{\bf\tilde{b}}\nabla_{x}\cdot F({\bf U})-\varepsilon\Delta t{\bf\tilde{b}}\nabla_{x}\cdot{\bf H}({\bf U}),$ (3.53)
where ${\bf{\tilde{b}}}:=\tilde{b}^{\top}\otimes_{K}I_{d_{v}+2}$ and
$\nabla_{x}\cdot{\bf H}({\bf U})$ $=(\nabla_{x}\cdot H(U^{(1)}),\nabla_{x}\cdot H(U^{(2)}),...,\nabla_{x}\cdot H(U^{(s)}))^{\top}\in\mathbb{R}^{(2+d_{v})s},$
$\nabla_{x}\cdot H(U^{(i)})$ $=\left(\begin{array}[]{c}0\\ \nabla_{x}\cdot(\rho^{(i)}\Theta_{1}^{(i)})\\ \nabla_{x}\cdot({\mathbb{Q}_{1}^{(i)}}+\rho^{(i)}{\Theta_{1}^{(i)}}{u^{(i)}})\end{array}\right)\in\mathbb{R}^{2+d_{v}}\,(i=1,...,s).$
Now we evaluate the quantities $\Theta_{1}^{(i)}$ and $\mathbb{Q}_{1}^{(i)}$, for i=1,,si=1,...,s. To achieve this, we expand the anisotropic Gaussian $\mathcal{G}$ with respect to $\varepsilon$, i.e.,
@@
$\frac{\mathcal{M}[{{\bf G}}]-\mathcal{\mathcal{M}}[f^{n}]{\bf e}}{\Delta t}+\tilde{A}\,\left(L(\mathcal{M}[{\bf G}])\right)=\tilde{A}{diag\left(\mathcal{M}[{\bf G}]\right)}\left(A({\textbf{V}}):\frac{\sigma({\bf u})}{2}+2B({\textbf{V}})\cdot\nabla_{x}\sqrt{{\bf T}}\right)+\mathcal{O}(\varepsilon),$ (3.57)
where $\bf G$} denotes the vector of internal stage for the equation (A.112) and
$\begin{split}{\bf u}&=(u^{(1)},...,u^{(s)})^{\top}\in\mathbb{R}^{d_{v}s},{\bf T}=(T^{(1)},...,T^{(s)})^{\top}\in\mathbb{R}^{s},\\ \textbf{V}&=(V^{(1)},V^{(2)},...,V^{(s)})^{\top}\in\mathbb{R}^{d_{v}s}\quad V^{(i)}=\frac{v-u^{(i)}}{\sqrt{T^{(i)}}}\in\mathbb{R}^{d_{v}},\\ &A(\textbf{V}):\frac{\sigma({\bf u})}{2}\in\mathbb{R}^{s},\quad\left(A(\textbf{V}):\frac{\sigma({\bf u})}{2}\right)_{i}=\left(V^{(i)}\otimes V^{(i)}-\frac{1}{d_{v}}|V^{(i)}|^{2}Id\right):\sigma(u^{(i)})\\ &B(\textbf{V})\cdot\nabla_{x}\sqrt{{\bf T}}\in\mathbb{R}^{s},\quad\left(B(\textbf{V})\cdot\nabla_{x}\sqrt{{\bf T}}\right)_{i}=\frac{1}{2}V^{(i)}(|V^{(i)}|^{2}-(d_{v}+2))\cdot\nabla_{x}\sqrt{{\bf T}^{(i)}},\quad i=1,...,s,\end{split}$
Integrating both sides of (3.57) macroscopic variables are associated to $\bf G$} are obtained as follows:
$\begin{split}{\bf U}_{\bf G}&={{\bf e}\otimes_{K}U^{n}}-\Delta t{\bf\tilde{A}}\nabla_{x}\cdot F({\bf U}_{\bf G})+\mathcal{O}(\varepsilon\Delta t),\end{split}$ (3.58)
where ${\bf U}_{\bf G}$ is the macroscopic quantities associated to the stage values. Since the first stage values of $\bf F$} and $\bf G$} are the same $f^{n}$, $\bf U$} and ${\bf U}_{G}$ are obtained from the same ${\bf U}^{n}$ and this implies that $\|{\bf U}-{\bf U}_{\bf G}\|_{\infty}=\mathcal{O}(\varepsilon\Delta t)$. Then, by Lemma A.1 we have $\mathcal{M}[{\bf G}]=\mathcal{M}[{\bf F}]+\mathcal{O}(\varepsilon\Delta t)$. Inserting this into (3.56), we get
$\begin{split}\bar{\tau}({\bf f}_{1}-{\bf g})&=-A^{-1}\left(\displaystyle\frac{\mathcal{M}[{\bf G}]-\mathcal{M}[f^{n}]{\bf e}}{\Delta t}+\tilde{A}\,L(\mathcal{M}[{\bf G}])+\mathcal{O}(\varepsilon)+\mathcal{O}(\varepsilon\Delta t)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)\\ &=-A^{-1}\left(\tilde{A}{diag\left(\mathcal{M}[{\bf G}]\right)}\left(A({\textbf{V}}):\frac{\sigma({\bf u})}{2}+2B({\textbf{V}})\cdot\nabla_{x}\sqrt{{\bf T}}\right)+\mathcal{O}(\varepsilon)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)+\mathcal{O}(\varepsilon\Delta t)\end{split}$ (3.59)
@@
$\bar{\tau}({\bf f}_{1}-{\bf g})$ $=-A^{-1}\left(\tilde{A}{diag\left(\mathcal{M}[{\bf F}]\right)}\left(A({\textbf{V}}):\frac{\sigma({\bf u})}{2}+2B({\textbf{V}})\cdot\nabla_{x}\sqrt{{\bf T}}\right)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)+\mathcal{O}(\varepsilon)+\mathcal{O}(\varepsilon\Delta t)$
@@
${\bf f}_{1}={\bf g}-{\bar{\tau}}^{-1}A^{-1}\left(\tilde{A}{diag\left(\mathcal{M}[{\bf F}]\right)}\left(A({\textbf{V}}):\frac{\sigma({\bf u})}{2}+2B({\textbf{V}})\cdot\nabla_{x}\sqrt{{\bf T}}\right)\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)+\mathcal{O}(\varepsilon).$ (3.60)
Multiplying by $v\otimes v$ or $v|v|^{2}/2$ both side in (3.60) and taking the integration over $v$, we get from (A.113), (A.114) and (A.115) that
$\begin{split}{\rho^{(i)}\Theta_{1}^{(i)}}&=\nu{\rho^{(i)}\Theta_{1}^{(i)}}-{{\frac{1}{\tau^{(i)}}\sum_{j=1}^{s}(A^{-1}\tilde{A})_{ij}}}\rho^{(j)}T^{(j)}\sigma(u^{(j)})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right),\quad i=1,...,s\\ \mathbb{Q}_{1}^{(i)}+\rho^{(i)}\Theta_{1}^{(i)}u^{(i)}&=\nu{\rho^{(i)}\Theta_{1}^{(i)}}u^{(i)}-{{\frac{1}{\tau^{(i)}}\sum_{j=1}^{s}(A^{-1}\tilde{A})_{ij}}}\left(\rho^{(j)}T^{(j)}\sigma(u^{(j)})u^{(j)}+\frac{d_{v}+2}{2}\rho^{(j)}T^{(j)}\nabla_{x}T^{(j)}\right)+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right).\end{split}$
Thus, the approximations of stress tensor and heat flux are given by
$\rho{\bf\Theta_{1}}$ $=\begin{pmatrix}\rho^{(1)}{\Theta_{1}^{(1)}},\rho^{(2)}{\Theta_{1}^{(2)}},\cdots,\rho^{(s)}{\Theta_{1}^{(s)}}\end{pmatrix}^{\top}$
$=-\frac{1}{1-\nu}\left({\bar{\tau}}^{-1}A^{-1}\tilde{A}\,diag(\bar{p})\otimes_{K}Id\right)\times\sigma({\bf u})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
$=-\left(\bar{\mu}\otimes_{K}Id\right)\times\sigma({\bf u})+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
${\bf q}=\mathbb{Q}_{1}$ $=\begin{pmatrix}\mathbb{Q}_{1}^{(1)},\mathbb{Q}_{1}^{(2)},\cdots,\mathbb{Q}_{1}^{(s)}\end{pmatrix}^{\top}$
$=-\frac{d_{v}+2}{2}\left({\bar{\tau}}^{-1}A^{-1}\tilde{A}\,diag(\bar{p})\otimes_{K}Id\right)\times\nabla_{x}{\bf T}+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
$=-\left(\bar{\kappa}\otimes_{K}Id\right)\times\nabla_{x}{\bf T}+\mathcal{O}\left(\frac{\varepsilon}{\Delta t}\right)$
where $\sigma(u^{(i)})=\nabla_{x}u^{(i)}+(\nabla_{x}u^{(i)})^{\top}-\frac{2}{d_{v}}\nabla_{x}\cdot u^{(i)}Id$, and $\bar{p}=(p^{(1)},p^{(2)},...,p^{(s)})^{\top}$ with $p^{(i)}=\rho^{(i)}T^{(i)}$. Here we defined the viscosity and thermal conductivity $s\times s$ matrices ${\bar{\mu}}=(\mu_{ij})$ and ${\bar{\kappa}}=(\kappa_{ij})$ as
${\bar{\mu}}=\frac{1}{(1-\nu)}{\bar{\tau}}^{-1}(A^{-1}\tilde{A})\,diag(\bar{p}),\quad{\bar{\kappa}}=\frac{d_{v}+2}{2}\,{\bar{\tau}}^{-1}(A^{-1}\tilde{A})\,diag(\bar{p}).$ (3.61)
This implies that
${\bar{\tau}}^{-1}(A^{-1}\tilde{A})\,diag(\bar{p})=(1-\nu){\bar{\mu}}=\frac{2}{d_{v}+2}{\bar{\kappa}}$
@@
This form is consistent to the Prandtl number of the ES-BGK model. To sum up, the scheme reads
${\bf U}={\bf e}\otimes_{K}U^{n}-\Delta t{\bf\tilde{A}}\nabla_{x}\cdot F({\bf U})+\varepsilon\Delta t{\bf\tilde{A}}\nabla_{x}\cdot\underline{S}({\bf U})+\mathcal{O}(\varepsilon^{2}),$ (3.62)
and
$U^{n+1}=U^{n}-\Delta t{\bf\tilde{b}}\nabla_{x}\cdot F({\bf U})+\varepsilon\Delta t{\bf\tilde{b}}\nabla_{x}\cdot\underline{S}({\bf U})+\mathcal{O}(\varepsilon^{2}),$ (3.63)
for the numerical solution, where
$\begin{split}\nabla_{x}\cdot\underline{S}({\bf U})&=\left(\nabla_{x}\cdot\underline{S}(U^{(1)}),\nabla_{x}\cdot\underline{S}(U^{(2)}),...,\nabla_{x}\cdot\underline{S}(U^{(s)})\right)^{\top}\in\mathbb{R}^{(d_{v}+2)s},\\ \nabla_{x}\cdot\underline{S}(U^{(i)})&=\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0},\\ \nabla_{x}\cdot\left(\sum_{j=1}^{i-1}\bar{\mu}_{ij}\sigma(u^{(j)})\right),\\ \nabla_{x}\cdot\left(\sum_{j=1}^{i-1}\bar{\mu}_{ij}\sigma(u^{(j)})u^{(j)}+{\bar{\kappa}_{ij}\nabla_{x}T^{(j)}}\right)\end{array}\right)\\ &=\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0},\\ \nabla_{x}\cdot\left(\frac{1}{\tau^{(i)}}\sum_{j=1}^{i-1}\frac{1}{(1-\nu)}(A^{-1}\tilde{A})_{ij}p^{(j)}\sigma(u^{(j)})\right),\\ \nabla_{x}\cdot\left(\frac{1}{\tau^{(i)}}\sum_{j=1}^{i-1}\frac{1}{(1-\nu)}(A^{-1}\tilde{A})_{ij}p^{(j)}\sigma(u^{(j)})u^{(j)}+{\frac{d_{v}+2}{2}(A^{-1}\tilde{A})_{ij}p^{(j)}\nabla_{x}T^{(j)}}\right)\end{array}\right)\in\mathbb{R}^{d_{v}+2}.\end{split}$
For each i,j=1,,si,j=1,...,s, we now consider the approximation $\tau^{(i)}=\tau^{(j)}+\mathcal{O}(\Delta t)+\mathcal{O}(\varepsilon^{2})$. Inserting this approximation into the $\underline{S}({\bf U})$ in (3.62) and (3.63), we obtain
$\begin{split}{\bf U}&={\bf e}\otimes_{K}U^{n}-\Delta t{\bf\tilde{A}}\nabla_{x}\cdot F({\bf U})+\varepsilon\Delta t{\bf\tilde{A}}\nabla_{x}\cdot\underline{\underline{S}}({\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2}),\\ U^{n+1}&=U^{n}-\Delta t{\bf\tilde{b}}\nabla_{x}\cdot F({\bf U})+\varepsilon\Delta t{\bf\tilde{b}}\nabla_{x}\cdot\underline{\underline{S}}({\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2}),\end{split}$
where
@@
with $\mu^{(j)}=\frac{1}{(1-\nu)\tau^{(j)}}p^{(j)}$, $\kappa^{(j)}=\frac{d_{v}+2}{2\tau^{(j)}}p^{(j)}$. Defining
$\begin{split}\nabla_{x}\cdot S({\bf U})&=\left(\nabla_{x}\cdot S(U^{(1)}),\nabla_{x}\cdot S(U^{(2)}),...,\nabla_{x}\cdot S(U^{(s)})\right)^{\top}\in\mathbb{R}^{(d_{v}+2)s},\\ \nabla_{x}\cdot S(U^{(i)})&=\left(\begin{array}[]{c}\nabla_{x}\cdot{\bf 0},\\ \nabla_{x}\cdot\left(\mu^{(i)}\sigma(u^{(i)})\right),\\ \nabla_{x}\cdot\left(\mu^{(i)}\sigma(u^{(i)})u^{(i)}+\kappa^{(i)}\nabla_{x}T^{(i)}\right)\end{array}\right)\in\mathbb{R}^{d_{v}+2},\end{split}$
we finally derive
$\begin{split}{\bf U}&={\bf e}\otimes_{K}U^{n}-\Delta t{\bf\tilde{A}}\nabla_{x}\cdot F({\bf U})+\varepsilon\Delta t{\bf B}\nabla_{x}\cdot S({\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2}),\\ U^{n+1}&=U^{n}-\Delta t{\bf\tilde{b}}\nabla_{x}\cdot F({\bf U})+\varepsilon\Delta t{\boldsymbol{\omega}}\nabla_{x}\cdot S({\bf U})+\mathcal{O}(\varepsilon^{2})+\mathcal{O}(\varepsilon\Delta t^{2}),\end{split}$ (3.64)
where ${\bf B}=(\tilde{A}A^{-1}\tilde{A})\otimes_{K}I_{d_{v}+2}$ and ${\boldsymbol{\omega}}=(\tilde{b}^{\top}A^{-1}\tilde{A})\otimes_{K}I_{d_{v}+2}$. Next, we consider the following explicit-type RK method of order $p$ based on the coefficients matrices $\tilde{A},B$ and the weights $\tilde{b},w$ with ${B}=\tilde{A}A^{-1}\tilde{A},\,\omega^{\top}=\tilde{b}^{\top}A^{-1}\tilde{A}$ applied to the CNS equations (1.21):
$\begin{split}{\bf U}&={\bf e}\otimes_{K}U^{n}-\Delta t{\bf\tilde{A}}\nabla_{x}\cdot F({\bf U})+\varepsilon\Delta t{\bf B}\nabla_{x}\cdot{S}({\bf U}),\\[5.69054pt] U^{n+1}&=U^{n}-\Delta t{\bf\tilde{b}}\nabla_{x}\cdot{\bf F}({\bf U})+\varepsilon\Delta t{\boldsymbol{\omega}}\nabla_{x}\cdot{S}({\bf U}).\end{split}$
Assuming that $u(t)$ is the true solution of the CNS, then the local truncation error is $\mathcal{O}(\Delta t^{p})$, i.e.,
$\begin{split}\mathcal{O}(\Delta t^{p})&=\frac{u(t_{n+1})-u(t_{n})}{\Delta t}+\sum_{i=1}^{s}\tilde{b}_{i}\nabla_{x}\cdot F(u^{(i)})-\sum_{i=1}^{s}\omega_{i}\varepsilon\nabla_{x}\cdot S(u^{(i)})\end{split}$ (3.65)
@@
${u}^{(i)}=u(t_{n})-\Delta t\sum_{j=1}^{{i-1}}{\tilde{a}_{ij}}\nabla_{x}\cdot F(u^{(j)})-\Delta t\sum_{j=1}^{{i-1}}{{B}_{ij}}\varepsilon\nabla_{x}\cdot S(u^{(j)}),$
where $B=(B_{ij})$. Now, we consider the local trunction of our method (3.64):
$\begin{split}L.T.E.&=\frac{u(t_{n+1})-u(t_{n})}{\Delta t}+\sum_{i=1}^{s}\tilde{b}_{i}\left(\nabla_{x}F(u^{(i)})\right)-\sum_{i=1}^{s}\omega_{i}\left(\varepsilon\nabla_{x}S(u^{(i)})\right)+\mathcal{O}\left(\frac{\varepsilon^{2}}{\Delta t}\right)+\mathcal{O}\left(\varepsilon\Delta t\right).\end{split}$ (3.66)
where $\omega=(\omega_{1},\omega_{2},...,\omega_{s})^{\top}$. This combined with (3.65), gives
$L.T.E.=\mathcal{O}(\Delta t^{p})+\mathcal{O}\left(\frac{\varepsilon^{2}}{\Delta t}\right)+\mathcal{O}\left(\varepsilon\Delta t\right).$ (3.67)
Theorem 3.6 implies that, to accurately capture the Navier-Stokes equation, it is necessary that the local truncation error satisfies $L.T.E=o(\varepsilon)$, which holds true if $\Delta t^{p}+\varepsilon\Delta t+\frac{\varepsilon^{2}}{\Delta t}={o}(\varepsilon)$.
@@
Setting $\tau=1$ in (1.7) the macroscopic explicit-type RK scheme has a local truncation error: $L.T.E.=\mathcal{O}(\Delta t^{p})+\mathcal{O}(\varepsilon^{2}/\Delta t).$
@@
$\bullet$ IMEX-RK scheme of type II. In this section we analyze the asymptotic behavior of type II IMEX-RK schemes. As type I scheme, at this stage, we do not need to assume that the scheme satisfies the GSA condition.
@@
For small values of $\varepsilon$ and with $\Delta t^{p}+\varepsilon\Delta t+\frac{\varepsilon^{2}}{\Delta t}={o}(\varepsilon)$, the IMEX-RK of type II (2.30) satisfying
$\hat{\tilde{A}}\hat{A}^{-1}a=\tilde{a},\quad\hat{\tilde{b}}^{\top}\hat{A}^{-1}a=\tilde{b}_{1},$ (3.68)